For fiscal year 2018, Walla Walla Department Store had net income of $6,010,000. Interest expense was $2,253,750, and the company’s tax rate on income was 40 percent. Total assets were $80,351,000, and noninterest-bearing current liabilities were $7,148,100. The company’s cost of capital (required rate of return) is 10 percent.

Calculate NOPAT, invested capital, and residual income for Walla Walla Department Store. (Enter negative answers preceding either - sign, e.g. -45 or in parentheses, e.g. (45).)

NOPAT $

Invested capital $

Residual income $


Comment on the company’s profitability.

The company appears

profitablenot profitable

.
Answer #1

NOPAT

= 6,010,000 + (2,253,750 - 40%)

= 7,362,250

Invested capital = 80,351,000 - 7,148,100

= 73,202,900

Residual income = NOPAT - Required return

= 7,362,250 - (73,202,900*10%)

= 41,960

Profitable

Comment if you face any issues